# Simulated Laser Cooling and Magneto-Optical Trapping of Group IV Atoms

### v1

@@ Abstract
We present a scheme for laser cooling and magneto-optical trapping of the Group IV (a.k.a. Group 14 or tetrel) atoms silicon (Si), germanium (Ge), tin (Sn), and lead (Pb). These elements each possess a strong Type-II transition ($J\rightarrow J^{\prime}=J-1$) between the metastable $s^{2}p^{2}\,{}^{3}P_{1}$ state and the excited $s^{2}ps^{\prime}\,{}^{3}P_{0}^{o}$ state at an accessible laser wavelength, making them amenable to laser cooling and trapping. We focus on the application of this scheme to Sn, which has several features that make it attractive for precision measurement applications. We perform numerical simulations of atomic beam slowing, capture into a magneto-optical trap (MOT), and subsequent sub-Doppler cooling and compression in a blue-detuned MOT of Sn atoms. We also discuss a realistic experimental setup for realizing a high phase-space density sample of Sn atoms.
@@ I Introduction
Adding a new class of elements to the list of laser-coolable species would constitute another leap in the quest to control and manipulate quantum matter. In this work, we present numerical simulations to verify a scheme for laser cooling and trapping of Group IV elements — silicon (Si), germanium (Ge), tin (Sn), and lead (Pb) — using a Type-II transition. Among these elements, previous attempts have been made to study pathways towards producing ultracold gases of Si atoms [40, 41, 42, 43], but a Si MOT has not been realized to date. Our work should provide a general path forward for laser cooling and trapping of Group IV elements for applications in quantum science and precision measurement.
@@ II.1 Structure of Group IV Atoms
In their ground state, all Group IV elements have four valence electrons in a $s^{2}p^{2}$ atomic orbital configuration. Generically, there are five even parity states in the ground-state manifold, with terms $^{3}P_{0,1,2},{{}^{1}D}_{2}$, and $^{1}S_{0}$ in order of energy from lowest to highest. The lowest-lying odd parity excited state has configuration $s^{2}ps^{\prime}$ and term $^{3}P_{0}^{o}$. [^1] C and Si also have a $sp^{3}$ $^{5}S_{2}^{o}$ state below the $s^{2}ps^{\prime}$ $^{3}P^{o}$ states and above the $s^{2}p^{2}$ $^{1}S_{0}$ state, but this does not affect our discussion. All elements in this group have multiple spin-zero ($I=0$) even isotopes and several spinful odd isotopes; in this paper we focus only on the even isotopes. A level diagram of this structure is shown in Fig. 1.
@@
We propose using the metastable $s^{2}p^{2}\,{{}^{3}}P_{1}$ to excited $s^{2}ps^{\prime}\,{{}^{3}}P_{0}^{o}$ transition in the Group IV elements as a pure Type-II optical cycling transition that is amenable to laser cooling and trapping. This transition is dipole-allowed with large natural linewidth. Moreover, it is completely closed by selection rules, meaning that no repumper laser is necessary. [^2] In Pb, the metastable state lifetime is short enough that a $^{3}P_{0}\rightarrow{{}^{3}}P_{1}$ repumper laser ($\lambda=1279$ nm) may be useful. Table 1 lists the relevant properties of this transition across all Group IV elements, as well as the metastable state lifetime. In the case of Si, Ge, Sn, and Pb, the transition wavelength can be produced with current ultraviolet (UV) laser technology. In the case of C, the transition wavelength is too short to currently be produced for laser cooling and trapping, so we do not consider it further in this work.
Table 1: Optical transition wavelength $\lambda_{c}$, natural linewidth $\Gamma$, and saturation intensity $I_{\text{sat}}$ of the $s^{2}p^{2}\,{{}^{3}}P_{1}\rightarrow s^{2}ps^{\prime}\,{{}^{3}}P_{0}^{o}$ transition proposed for use in laser cooling and trapping of Group IV elements [44]. The metastable $s^{2}p^{2}\,{{}^{3}}P_{1}$ state lifetime $\tau_{g}$ is also listed [44]. More details on calculating $\tau_{g}$ are given in the Supplemental Material (SM) [45].
Element $\lambda_{c}$ (nm) $\Gamma$ (MHz) $I_{\text{sat}}$ (mW/cm2) $\tau_{g}$ (s)
C 166 $2\pi\times 55$ 1572 $1.3\times 10^{7}$
Si 252 $2\pi\times 35$ 286 $1.2\times 10^{5}$
Ge 271 $2\pi\times 45$ 296 $3.2\times 10^{2}$
Sn 303 $2\pi\times 32$ 149 11
Pb 368 $2\pi\times 22$ 58 0.14
We note that a more conventional-seeming optical cycling transition is also present: $s^{2}p^{2}\,{}^{3}P_{2}$ to $sp^{3}\,{{}^{3}}D_{3}^{o}$ (for Si) or $s^{2}pd\,{{}^{3}}D_{3}^{o}$ (for Ge and Sn), i.e. the excited state is the lowest-lying odd-parity $D$ state. This was used to attempt laser cooling of Si ($\lambda=222\,\text{nm}$) [42]. This cycling transition is leaky, with an estimated branching ratio $^{3}D_{3}^{o}\rightarrow{{}^{1}}D_{2}$ of $\sim\!0.02\%-50\%$ depending on the element [44], necessitating a repumper laser to close the cycling transition. However, inclusion of this repumper creates a $\Lambda$-system, effectively making a Type-II cycling transition with two UV lasers on two different transitions. The cycling transition we propose is simpler and should lead to more effective laser cooling and trapping of Group IV elements.
In addition to supporting a Type-II transition, all Group IV elements possess four linearly independent clock transitions between the long-lived $s^{2}p^{2}$ states. Details on these transitions, which can only be of magnetic dipole (M1) or electric quadrupole (E2) type, are listed in Table 2.
@@
State Transition Type $\lambda$ (nm) $\Gamma_{\rm p}$ (Hz) $\Gamma_{\rm tot}$ (Hz)
$^{3}P_{1}$ $\leftrightarrow\,^{3}P_{0}$ M1 5911 $2\pi\times 0.01$ $2\pi\times 0.01$
$^{3}P_{2}$ $\leftrightarrow\,^{3}P_{0}$ E2 2917 $2\pi\times 10^{-4}$ $2\pi\times 0.01$
$\leftrightarrow\,^{3}P_{1}$ M1/E2 5761 $2\pi\times 0.01$
$^{1}D_{2}$ $\leftrightarrow\,^{3}P_{0}$   E2* 1161 $2\pi\times 10^{-5}$ $2\pi\times 0.16$
$\leftrightarrow\,^{3}P_{1}$   M1/E2* 1445 $2\pi\times 0.07$
$\leftrightarrow\,^{3}P_{2}$   M1/E2* 1929 $2\pi\times 0.08$
$^{1}S_{0}$ $\leftrightarrow\,^{3}P_{1}$   M1* 646 $2\pi\times 1.13$ $2\pi\times 1.42$
$\leftrightarrow\,^{3}P_{2}$   E2* 728 $2\pi\times 0.11$
$\leftrightarrow\,^{1}D_{2}$ E2 1170 $2\pi\times 0.18$
II.2 Applications of Laser-Cooled Group IV Atoms
Ultracold atomic gases of Group IV elements have potential applications ranging from improved quantum sensing and novel quantum computing platforms to precision tests of fundamental physics and searches for physics beyond the Standard Model (BSM). Initial efforts to produce an ultracold atomic gas of Si atoms focused on envisioned applications to solid-state-based nuclear spin quantum computing [47, 48] and quantum sensing [49]. Resonant photoionization of ultracold radioactive $^{31}\text{Si}$ atoms in a MOT, followed by ion-guiding optics, would enable nm-scale precision implantation of $^{31}\text{Si}^{+}$ ions in a Si substrate [42, 50]. Subsequent $\beta$-decay of these ions ($\tau\sim\!3$ hrs) would create an array of $^{31}\text{P}^{+}$ ions doped in the Si substrate, realizing the building blocks of the Kane quantum computing architecture [47].
The plethora of accessible clock transitions in the ground $s^{2}p^{2}$ configuration can be utilized for a variety of precision measurement experiments. Sn is particularly attractive for such applications because it possesses a chain of seven stable, $I=0$ isotopes (the longest found in nature). The combination of many clock transitions and many $I=0$ isotopes makes laser-cooled and trapped Sn atoms a promising platform for probing non-linearity in King plot isotope shift measurements [51, 52]. Detection of non-linearity could indicate BSM physics manifesting as a new, light boson that interacts with electrons and nucleons [53]. Other charge states of Sn, such as $\text{Sn}^{2+}$ and $\text{Sn}^{7+}\text{-}\text{Sn}^{10+}$ ions, have been recently explored and also possess suitable clock transitions for such measurements [54, 55, 56, 57].
@@ II.2 Applications of Laser-Cooled Group IV Atoms
II.3 Scheme for Laser Cooling and Trapping with the $^{3}P_{1}\rightarrow{{}^{3}P}_{0}^{o}$ Transition
@@
In this work, we assume the source to have the same properties as the CBGB used in SrF laser cooling and trapping [84]. The buffer gas is helium at $T\sim\!4$ K, with flow rate $\mathcal{F}\sim\!5$ sccm. The initial longitudinal velocity distribution is roughly Gaussian, with mean longitudinal velocity $\bar{v}_{L}=140\,\text{m/s}$ and full width at half maximum (FWHM) $\Delta{v_{L}}=75\,\text{m/s}$. The initial transverse velocity distribution is also roughly Gaussian, with mean $\bar{v}_{T}=0\,\text{m/s}$ and FWHM $\Delta{v_{T}}=75\,\text{m/s}$. We assume a beam brightness across all isotopes of $\mathcal{B}\sim\!1.5\times 10^{12}$ atoms in the $^{3}P_{1}$ metastable state/steradian/pulse, commensurate with the measured brightness of Sr atoms in the SrF beam [85] after accounting for the metastable state population. We remark that using a thermal beam source of atoms may also be viable, but we do not explore this path here.
@@ III Numerical Simulations of Optical Forces
From $a(r,v)$, we can evolve atomic trajectories in phase space. These trajectories incorporate stochastic photon recoils from spontaneous emission as atoms repeatedly absorb and emit photons from the cooling laser. Further details of our numerical simulation analysis are given in the Supplementary Material (SM) [45].
@@ IV Capture into Red-Detuned MOT
Figure 2: Dual frequency/polarization mechanism used for capture and conveyor belt MOTs of Sn atoms. Both configurations consist of two co-propagating laser beams with orthogonal circular polarizations. $\Delta$ is the detuning of a laser frequency from the $^{3}P_{1}\rightarrow{{}^{3}}P_{0}^{o}$ resonance.
The optimized and experimentally viable parameters for this stage are shown in Table 3. With Gaussian laser beams of $w_{\text{MOT}}=7$ mm ($1/e^{2}$ intensity radius), the scheme requires a power $P\sim\!200\,\text{mW}$ per beam of the MOT. This can be achieved with a total of $P\approx 200\,\text{mW}$ by using a single-beam configuration as in Ref. [36, 86]. The axial magnetic (B-)field gradient is $\partial B/\partial z=50\,\text{G/cm}$.
Table 3: Parameters used for Sn capture and compressed MOT simulations (see Fig. 3 and Fig. 6). For the compression stage, the parameters shown are those for the last stage of compression. Laser power $P_{i}$ and detuning $\Delta_{i}$ are given for each polarization component $\hat{p}_{i}$ in the dual frequency/polarization scheme. $\partial B/\partial z$ is the axial B-field gradient. The Gaussian laser beam size ($1/e^{2}$ intensity radius) is $w_{\text{MOT}}=7$ $m$ in both MOT stages. In the results, $v_{\rm cap}$ is the capture velocity; $T$ and $\sigma$ are the temperature and cloud size; $P_{\rm exc}$ is the average population in the excited state.
MOT Stage $\partial B/\partial z$ (G/cm) Laser Parameters Results
$P_{i}$ (mW) $\Delta_{i}$ ($\Gamma$) $\hat{p}_{i}$ $v_{\rm cap}$ (m/s) $T$ (mK) $\sigma$ ($m$) $P_{\rm exc}$ (%)
Capture 50 172 1.0-1.0 $\sigma^{-}$ 28.5 225 1.9 7.8
34 +0.5+0.5 $\sigma^{+}$
Compress 150 5 0.5-0.5 $\sigma^{-}$ 8 0.5 0.6
1 +0.25+0.25 $\sigma^{+}$
A map of the acceleration experienced by Sn atoms in the capture MOT is shown in Fig. 3. The plot clearly indicates the presence of strong Doppler cooling at large velocities, for example with peak deceleration $a\approx-40$ mm/ms2 at $v\approx 20$ m/s and $r\approx-4$ mm. At lower velocities, $v\lesssim 5\,\text{m/s}$, the acceleration changes sign, indicating the presence of sub-Doppler heating. This behavior is typical for Type-II transitions [88], and can be more clearly seen in plots of the spatial and velocity-averaged forces derived from integrating $a(r,v)$ over $v$ and $r$, respectively, as shown in the SM  [45].
@@
To determine capture velocity, we initialize atoms at r=14r=-14 mm, with increasing velocities $v$. An atom is deemed captured if, after $t=50$ ms of evolution, its trajectory has converged onto a steady-state oscillation in phase space, with the spatial extent of that oscillation $r_{\rm max}<w_{\rm MOT}$. We find that this condition is met for all trajectories for which the velocity reverses sign before the displacement along $\hat{r}$ exceeds $w_{\rm{MOT}}=7$ mm. We find a capture velocity of $v_{c}\approx 28.5$ m/s. This value is 2-3 times higher than what is experimentally measured for laser-cooled molecules like CaF and SrF [79, 89], due to the larger natural linewidth of the cooling transition and $\sim\!2\times$ higher photon momentum in Sn. Plots of trajectories for determining $v_{c}$ are shown in the SM [45].
To characterize the temperature and size of the MOT, we extend the trajectory of captured and trapped atoms by another $t=50$ ms to allow MOT observables to reach equilibrium. Using these trajectories, we calculate the rms displacement $\sigma$ (which we denote the cloud size), velocity $v_{\sigma}$, and temperature $T=mv_{\sigma}^{2}/k_{B}$, by averaging the position and velocity over the last 10 ms. The final size and temperature are averaged over ten such trajectories. For the capture MOT stage, we find $T\sim\!225\,\text{mK}$ and $\sigma\sim\!1.9\,\text{mm}$. For context, the Doppler temperature for this cooling transition in Sn is $T_{D}=760\,\mu\text{K}$. We attribute the extremely hot and large size of the capture red MOT to the large natural linewidth and Type-II nature of the optical cycling transition. A typical in-MOT trajectory is shown in the SM [45].
We note the possibility of two-photon ionization (TPI) when using the $^{3}P_{1}\rightarrow{{}^{3}}P_{0}^{o}$ optical cycling transition in Sn. This would manifest as one-body loss of the form $dN/dt=-\alpha N$. The loss rate due to TPI, $\alpha$, is given by:

### v2

@@ Abstract
We present a scheme for laser cooling and magneto-optical trapping of the Group IV (a.k.a. Group 14 or tetrel) atoms silicon (Si), germanium (Ge), tin (Sn), and lead (Pb). These elements each possess a strong Type-II transition ($J\rightarrow J^{\prime}=J-1$) between the metastable $s^{2}p^{2}\,{}^{3}\!P_{1}$ state and the excited $s^{2}ps^{\prime}\,{}^{3}\!P_{0}^{\circ}$ state at an accessible laser wavelength, making them amenable to laser cooling and trapping. We focus on the application of this scheme to Sn, which has several features that make it attractive for precision measurement applications. We perform numerical simulations of atomic beam slowing, capture into a magneto-optical trap (MOT), and subsequent sub-Doppler cooling and compression in a blue-detuned MOT of Sn atoms. We also discuss a realistic experimental setup for realizing a high phase-space density sample of Sn atoms.
@@ I Introduction
Adding a new class of elements to the list of laser-coolable species would constitute another leap in the quest to control and manipulate quantum matter. In this work, we present numerical simulations to verify a scheme for laser cooling and trapping of nearly all [^1] The vacuum UV transition wavelength for carbon (C) cannot currently be produced with sufficient continuous-wave power for laser cooling and trapping (see Table 1), so we do not consider it here. Group IV elements — silicon (Si), germanium (Ge), tin (Sn), and lead (Pb) — using a Type-II transition. Among these elements, previous attempts have been made to study pathways towards producing ultracold gases of Si atoms [53, 32, 82, 87], but a Si MOT has not been realized to date. Our work should provide a general path forward for laser cooling and trapping of Group IV elements for applications in quantum science and precision measurement.
@@ II.1 Structure of Group IV Atoms
In their ground state, all Group IV elements have four valence electrons in a $s^{2}p^{2}$ atomic orbital configuration. This leads to five even parity states in the ground-state manifold, with terms $^{3}\!P_{0,1,2},{{}^{1}\!D}_{2}$, and $^{1}\!S_{0}$ in order of energy from lowest to highest. The four excited states in this manifold cannot decay via electric dipole (E1) radiation and hence are all long-lived. The lowest-lying odd parity excited state has configuration $s^{2}ps^{\prime}$ and term $^{3}\!P_{0}^{\circ}$. [^2] C and Si also have a $sp^{3}$ $^{5}\!S_{2}^{\circ}$ state below the $s^{2}ps^{\prime}$ $^{3}\!P^{\circ}$ states and above the $s^{2}p^{2}$ $^{1}\!S_{0}$ state, but this does not affect our discussion. A level diagram of this structure is shown in Fig. 1.
@@
We propose using the metastable $s^{2}p^{2}\,{{}^{3}}\!P_{1}$ to excited $s^{2}ps^{\prime}\,{{}^{3}}\!P_{0}^{\circ}$ transition in the Group IV elements as a pure Type-II optical cycling transition that is amenable to laser cooling and trapping. This transition is dipole-allowed with large natural linewidth. Moreover, it is completely closed by selection rules, meaning that no repumper laser is necessary. [^3] In Pb, the metastable state lifetime is short enough that a $^{3}\!P_{0}\rightarrow{{}^{3}}\!P_{1}$ repumper laser ($\lambda=1279$ nm) may be useful. Table 1 lists the relevant properties of this transition across all Group IV elements, as well as the metastable state lifetime. In the case of Si, Ge, Sn, and Pb, the transition wavelength can be produced with sufficient power using current UV laser technology.
Table 1: Optical transition wavelength $\lambda_{c}$, natural linewidth $\Gamma$, and saturation intensity $I_{\text{sat}}$ of the $s^{2}p^{2}\,{{}^{3}}\!P_{1}\rightarrow s^{2}ps^{\prime}\,{{}^{3}}\!P_{0}^{\circ}$ transition proposed for use in laser cooling and trapping of Group IV elements [52]. The metastable $s^{2}p^{2}\,{{}^{3}}\!P_{1}$ state lifetime $\tau_{g}$ is also listed [52]. Details on calculating $\tau_{g}$ are given in Appendix A.
Element $\lambda_{c}$ (nm) $\Gamma/2\pi$ (MHz) $I_{\text{sat}}$ (mW/cm2) $\tau_{g}$ (s)
C 166 $55$ 1600 $1\times 10^{7}$
Si 252 $35$ 290 $1\times 10^{5}\!$
Ge 271 $45$ 300 $3\times 10^{2}$
Sn 303 $32$ 150 11
Pb 368 $22$ 60 0.14
II.2 Applications of Laser-Cooled Group IV Atoms
Ultracold atomic gases of Group IV elements have potential applications ranging from improved quantum sensing and novel quantum computing platforms to precision tests of fundamental physics and searches for physics beyond the Standard Model (BSM). Initial efforts to produce an ultracold atomic gas of Si atoms focused on envisioned applications to solid-state-based nuclear spin quantum computing [50, 54] and quantum sensing [68]. Resonant photoionization of ultracold radioactive $^{31}\text{Si}$ atoms in a MOT, followed by ion-guiding optics, would enable nm-scale precision implantation of $^{31}\text{Si}^{+}$ ions in a Si substrate [82, 41]. Subsequent $\beta$-decay of these nuclei ($\tau\sim\!3$ hrs) would create an array of $^{31}\text{P}^{+}$ ions doped in the Si substrate, realizing the building blocks of the Kane quantum computing architecture [50].
In addition to supporting a Type-II cycling transition, all Group IV elements possess various clock transitions that can address all four linearly independent energy differences between the long-lived $s^{2}p^{2}$ states. Details on these transitions, which can only be of magnetic dipole (M1) or electric quadrupole (E2) type, are listed in Table 2 for the case of Sn.
@@ II.2 Applications of Laser-Cooled Group IV Atoms
State Transition Type $\lambda$ (nm) $\Gamma_{\text{p}}/2\pi$ (Hz) $\Gamma_{\text{tot}}/2\pi$ (Hz)
$^{3}\!P_{1}$ $\leftrightarrow\,^{3}\!P_{0}$ M1 5911 $0.01$ $0.01$
$^{3}\!P_{2}$ $\leftrightarrow\,^{3}\!P_{0}$ E2 2917 $10^{-4}$ $0.01$
$\leftrightarrow\,^{3}\!P_{1}$ M1/E2 5761 $0.01$
$^{1}\!D_{2}$ $\leftrightarrow\,^{3}\!P_{0}$   E2* 1161 $10^{-5}$ $0.16$
$\leftrightarrow\,^{3}\!P_{1}$   M1/E2* 1445 $0.07$
$\leftrightarrow\,^{3}\!P_{2}$   M1/E2* 1929 $0.08$
$^{1}\!S_{0}$ $\leftrightarrow\,^{3}\!P_{1}$   M1* 646 $1.13$ $1.4$
$\leftrightarrow\,^{3}\!P_{2}$   E2* 728 $0.11$
$\leftrightarrow\,^{1}\!D_{2}$ E2 1170 $0.18$
The plethora of metastable states in the ground $s^{2}p^{2}$ configuration provides several accessible clock transitions that can be utilized for a variety of precision measurement experiments. Sn is particularly attractive for such applications because it possesses a chain of seven stable $I=0$ isotopes (the longest found in nature). The combination of many clock transitions and many $I=0$ isotopes makes laser-cooled and trapped Sn atoms a promising platform for probing non-linearity in King plot isotope shift measurements [9, 36]. Detection of non-linearity could indicate BSM physics manifesting as a new, light boson that interacts with electrons and nucleons [10]. Other charge states of Sn, such as $\text{Sn}^{2+}$ and $\text{Sn}^{7+}\text{-}\text{Sn}^{10+}$ ions, have been recently explored and also possess suitable clock transitions for such measurements [60, 95, 104, 64].
@@
II.3 Scheme for Laser Cooling and Trapping with the $^{3}\!P_{1}\rightarrow{{}^{3}\!P}_{0}^{\circ}$ Transition
@@
We note, in passing, that a more conventional-seeming optical cycling transition is also present: $s^{2}p^{2}\,{}^{3}\!P_{2}\leftrightarrow sp^{3}\,{{}^{3}}\!D_{3}^{\circ}$ (for Si) or $s^{2}pd\,{{}^{3}}\!D_{3}^{\circ}$ (for Ge and Sn), i.e. the excited state is the lowest-lying odd-parity triplet state with $J=3$. This was used to attempt laser cooling of Si ($\lambda=222\,\text{nm}$) [82]. This cycling transition is leaky, with an estimated branching ratio $^{3}\!D_{3}^{\circ}\rightarrow s^{2}p^{2}{{}^{1}}\!D_{2}$ of $\sim\!0.02\%-50\%$ depending on the element [52], necessitating a repumper laser to close the cycling transition. However, including this repumper creates a $\Lambda$-system, effectively making a Type-II cycling transition with two UV lasers on two different transitions. Our approach is simpler and should lead to more effective laser cooling and trapping of Group IV elements.
@@
In this work, we assume the source to have the same properties as the CBGB used in SrF laser cooling and trapping [8]. The buffer gas is helium at $T\sim\!4$ K, with flow rate $\mathcal{F}\sim\!5$ sccm. The initial longitudinal velocity distribution is roughly Gaussian, with mean longitudinal velocity $\bar{v}_{L}=140\,\text{m/s}$ and full width at half maximum (FWHM) $\Delta{v_{L}}=75\,\text{m/s}$. The initial transverse velocity distribution is also roughly Gaussian, with mean $\bar{v}_{T}=0\,\text{m/s}$ and FWHM $\Delta{v_{T}}=75\,\text{m/s}$. We assume a beam brightness across all isotopes of $\mathcal{B}\sim\!1.5\times 10^{12}$ atoms in the $^{3}\!P_{1}$ metastable state per steradian per pulse, commensurate with the measured brightness of Sr atoms in the SrF beam [49] after accounting for the metastable state population. We remark that using a thermal beam source of atoms may also be viable, but we do not explore this path here.
@@ III Numerical Simulations of Optical Forces
From $a(r,v)$, we can evolve atomic trajectories in phase space. These trajectories incorporate stochastic photon recoils from spontaneous emission as atoms repeatedly absorb and emit photons from the cooling laser. Further details of our numerical simulation analysis are given in Appendix B.
@@ IV Capture into Red-Detuned MOT
Figure 2: Dual frequency/polarization mechanism used for capture and conveyor belt MOTs of Sn atoms. Both configurations consist of two co-propagating laser beams with orthogonal circular polarizations. $\Delta$ is the detuning of a laser frequency from the $^{3}\!P_{1}\rightarrow{{}^{3}}\!P_{0}^{\circ}$ resonance.
The optimized and experimentally viable parameters for this stage are shown in Table 3. With Gaussian laser beams of $w_{\text{MOT}}=7$ mm ($1/e^{2}$ intensity radius), the scheme requires a power $P_{\text{beam}}\sim\!200\,\text{mW}$ per beam of the MOT. This can be achieved with a total of $P_{\text{tot}}\approx 200\,\text{mW}$ by using a single-beam configuration as in Ref. [56, 48]. The axial magnetic (B-)field gradient is $\partial B/\partial z=50\,\text{G/cm}$.
Table 3: Parameters used for Sn capture and compressed MOT simulations (see Fig. 3 and Fig. 6). For the compression stage, the parameters shown are those for the last stage of compression. Laser power $P_{i}$, saturation parameter $s_{i}$, and detuning $\Delta_{i}$ are given for each polarization component $\hat{p}_{i}$ (for the given polarization, we assume the beam wave vector $\hat{k}_{i}$ points towards $\hat{x}$ and that the B-field increases along $\hat{x}$) in the dual frequency/polarization scheme. $\partial B/\partial z$ is the axial B-field gradient. The Gaussian laser beam size ($1/e^{2}$ intensity radius) is $w_{\text{MOT}}=7$ $m$ in both MOT stages. In the results, $v_{\rm cap}$ is the capture velocity; $T$ and $\sigma$ are the temperature and cloud size; $P_{\rm exc}$ is the average population in the excited state.
MOT Stage $\frac{\partial B}{\partial z}$ (G/cm) Laser Parameters Results
$P_{i}$ (mW) $s_{i}$ $\Delta_{i}$ ($\Gamma$) $\hat{p}_{i}$ $v_{\rm cap}$ (m/s) $T$ (mK) $\sigma$ ($m$) $P_{\rm exc}$ (%)
Capture 50 172 1.5 1.0-1.0 $\sigma^{-}$ 28.5 225 1.9 7.8
34 0.3 +0.5+0.5 $\sigma^{+}$
Compress 150 5 0.05 0.5-0.5 $\sigma^{-}$ 8 0.5 0.6
1 0.01 +0.25+0.25 $\sigma^{+}$
A map of the acceleration experienced by Sn atoms in the capture MOT is shown in Fig. 3. The plot clearly indicates the presence of strong Doppler cooling at large velocities, for example with peak deceleration $a\approx-40$ mm/ms2 at $v\approx 20$ m/s and $r\approx-4$ mm. At lower velocities, $v\lesssim 5\,\text{m/s}$, the acceleration changes sign, indicating the presence of sub-Doppler heating. This behavior is typical for Type-II transitions [27], and can be more clearly seen in plots of the spatial and velocity-averaged forces derived from integrating $a(r,v)$ over $v$ and $r$, respectively, as shown in Appendix B.
@@
To determine capture velocity, we initialize atoms at r=14r=-14 mm, with increasing velocities $v$. An atom is deemed captured if, after $t=50$ ms of evolution, its trajectory has converged onto a steady-state oscillation in phase space, with the spatial extent of that oscillation $r_{\rm max}<w_{\rm MOT}$. We find that this condition is met for all trajectories for which the velocity reverses sign before the displacement along $\hat{r}$ exceeds $w_{\rm{MOT}}=7$ mm. We find a capture velocity of $v_{c}\approx 28.5$ m/s. This value is 2-3 times higher than what is experimentally measured for laser-cooled molecules like CaF and SrF [101, 89], due to the larger natural linewidth of the cooling transition and $\sim\!2\times$ higher photon momentum in Sn. Plots of trajectories for determining $v_{c}$ are shown in Appendix B.
To characterize the temperature and size of the MOT, we extend the trajectory of captured and trapped atoms by another $t=50$ ms to allow MOT observables to reach equilibrium. Using these trajectories, we calculate the rms displacement $\sigma$ (which we denote the cloud size), velocity $v_{\sigma}$, and temperature $T=mv_{\sigma}^{2}/k_{B}$, by averaging the position and velocity over the last 10 ms. The final size and temperature are averaged over ten such trajectories. For the capture MOT stage, we find $T\sim\!225\,\text{mK}$ and $\sigma\sim\!1.9\,\text{mm}$. For context, the Doppler temperature for this cooling transition in Sn is $T_{D}=760\,\mu\text{K}$. We attribute the extremely hot temperature and large size of the capture red MOT to the large natural linewidth and Type-II nature of the optical cycling transition; a similarly hot MOT was observed in Ref. [71]. A typical in-MOT trajectory is shown in Appendix B.
We note the possibility of two-photon ionization (TPI) when using the $^{3}\!P_{1}\rightarrow{{}^{3}}\!P_{0}^{\circ}$ optical cycling transition in Sn. This would manifest as one-body loss of the form $dN/dt=-\alpha N$. The loss rate due to TPI, $\alpha$, is given by: